# Gold after Randomized Sand: Model-X Split Knockoffs for Controlled Transformation Selection

### v1

@@ 2.2.2 Construction for Pairwise Comparisons
In this section, we discuss the case of pairwise comparisons where the distribution for the design matrix $X$ can not be approximated by Gaussian distributions.
For example, let a collection of pairwise comparison be represented by $\mathbf{X}\in\mathbb{R}^{n\times p},\mathbf{Y}\in\mathbb{R}^{n}$ where each row $X\in\mathbb{R}^{p}$ contains a random paired comparison such that $1$ denotes the first term in the random pair and $-1$ denotes the second term in the random pair and other values are 0. For example, $X=(1,-1,0,\dots,0)$ means that the comparison is done between the first and second objects. Each element $Y\in\mathbb{R}$ of $\mathbf{Y}$ is modeled from
@@
where $\beta^{*}\in\mathbb{R}^{p}$ represents the unknown true scores for $p$ objects, and $f(x)$ is a cumulative distribution function which is symmetrical at 0, i.e. it satisfies $f(x)=1-f(-x)$. In order to conduct pairwise comparisons, let $D$ be the graph difference operator of a fully connected graph. [^1] The construction detailed in this section is applicable when $D$ contains all the edges present in $\mathbf{X}$. Should this condition not be met, one can simply augment $D$ to include all required edges. Then the pairs with large differences will be selected in the model (1).
Sequential Constructions
One direct way to construct the pair $(A,\widetilde{A})$ is the sequential construction following Algorithm 1 with certain choice of $A$. The detailed construction for the above pairwise comparisons setting is shown in Supplementary Material Section F. However, as will be demonstrated later in Section 4.2, these constructions not only incur high computational costs but also exhibit reduced selection power due to the inherent dependency among $(\widetilde{X}^{A},A,\widetilde{A})$. This motivates our search for a simultaneous approach to construct Split Knockoff copies. Such an approach would not only reduce computational expense but also enhance selection power by ensuring $\widetilde{X}^{A}$ is independent of $A$, a characteristic previously demonstrated to be advantageous for selection power in the context of normal distributions.
Bootstrap+ Constructions
In this section, we present a novel method for generating Split Knockoff copies, which we term the bootstrap+ construction. This approach simultaneously constructs the pair $(A,\widetilde{A})$ and incorporates bootstrap sampling of the design matrix $\mathbf{X}^{r}$ after augmenting it with rows of $0_{p}$.
This construction initiates with an augmented bootstrap resampling procedure. Each row $(X^{r},Y^{r})$ of the resulting augmented and resampled design matrix $\mathbf{X}^{r}$ and response vector $\mathbf{Y}^{r}$ is generated from the distribution of $X$ by:
@@
    $\mathbb{P}(X^{r}=0_{p})=\mathbb{P}(X^{r}\neq 0_{p})=\frac{1}{2}$.  [^2] One can not use a general probability $\mathbb{P}(X^{r}=0_{p})=p_{0}\neq 1/2$ here since $p_{0}\neq 1/2$ will destroy the exchangeability of $\tilde{A}$ and $A$.
@@
    When $X^{r}\neq 0_{p}$, $X^{r}$ takes a random row from $\mathbf{X}$, and $Y^{r}$ takes the corresponding row in $\mathbf{Y}$.
This resampling enables the following construction of the Split Knockoff copies based on $(X^{r},Y^{r})$ at a much lower computational cost.
@@
    The random vector $A\in\mathbb{R}^{m}$ is generated to satisfy $\widetilde{X}^{A}_{r}:=X^{r}-AD=0_{p}$, i.e. if $X^{r}=0$ then $A=0$; if $X^{r}$ takes a random row of $\mathbf{X}$ then $A$ selects the same row in $D$. In this way, $\widetilde{X}^{A}_{r}$ is clearly independent from $A$, which should help improve the selection power as shown in the previous Gaussian case of $X\sim\mathcal{N}(0,\Sigma_{X})$.
@@
    The random vector $\widetilde{A}$ is generated based on $A$ by
    $\mathbb{P}(\widetilde{A}=0_{m}|A\neq 0_{m})=1,$
    $\mathbb{P}(\widetilde{A}=e_{i}|A=0_{m})=\mathbb{P}(A=e_{i}|A\neq 0_{m}),$
    for any canonical basis vector $e_{i}$.
With the above construction, one can easily verify that $\widetilde{X}^{A}_{r}\equiv 0_{p}$, while the vector $[A,\widetilde{A}]\in\mathbb{R}^{2m}$ is the one-hot vector with one element being 1 and other elements being zeros. Moreover, $\mathbb{P}[A_{j}=1]=\mathbb{P}[\widetilde{A}_{j}=1]$ for all $j$, which directly leads to the exchangeability.
This procedure clearly lets $\widetilde{X}^{A}_{r}$ independent from $A$, which should help improve the selection power as demonstrated in the Gaussian case. Simulation experiments validate the selection power improvement compared with the sequential construction in Section 4.2. Moreover, the bootstrap+ construction only requires $\mathbb{P}(A=e_{i}|A\neq 0_{m})$, which can be directly computed from the known empirical distribution $P_{\mathbf{X}}$ of the random vector $X$. This helps lower the computational cost, while avoiding the theoretical FDR inflation issues introduced from the estimation error on the distribution of $X$ at the same time.
@@ 2.3 Computation of $W$-statistics
With the above construction, we are now ready to compute the routine $W$-statistics similarly as in (Barber and Candès, 2015, Candes et al., 2018) through
$W_{i}=w_{i}\left\{(\widetilde{X}^{A},A,\widetilde{A}),Y\right\},$ (32)
for $i\in\{1,2,\cdots,m\}$. Similarly as in (Barber and Candès, 2015, Candes et al., 2018), the $W$-statistics needs to satisfy the flip-sign property on $A$ and $\widetilde{A}$,
$w_{j}\left\{(\widetilde{X}^{A},A,\widetilde{A})_{\text{swap}(S)},Y\right\}=\begin{cases}w_{j}\left\{(\widetilde{X}^{A},A,\widetilde{A})_{\text{swap}(S)},Y\right\},&j\notin S,\\ -w_{j}\left\{(\widetilde{X}^{A},A,\widetilde{A})_{\text{swap}(S)},Y\right\},&j\in S.\end{cases}$ (33)
Here, $S$ represents any subset of $\mathcal{H}^{\gamma}_{0}$, and $(\widetilde{X}^{A},A,\widetilde{A})_{\text{swap}(S)}$ denotes the operation of swapping the columns of $A$ and $\widetilde{A}$, while keeping $\widetilde{X}^{A}$ unchanged.
There are many approaches as shown in (Barber and Candès, 2015, Candes et al., 2018) in constructing such $W$-statistics satisfying the above property. In this paper, we adopt the following specific construction proposed by (Cao et al., 2024a) that searches in the relaxed neighborhood of the subspace constraint to optimize the selection power. Consider the following optimization problem for hyper-parameters $\lambda>0$, $\nu>0$:
@@
where the loss function $L$ is related with the model or link function, e.g., mean squared loss for linear regression, and cross-entropy loss for logistic regression. The $W$-statistics is constructed as
$W=|\gamma(\lambda)|-|\widetilde{\gamma}(\lambda)|.$ (35)
The hyper-parameters $\lambda>0$ and $\nu>0$ can be optimized to maximize the selection power. Here, $\lambda>0$ is the well-known $\ell_{1}$ penalization parameter. Meanwhile, $\nu>0$ is the parameter which controls the Euclidean relaxation gap between $D\beta$ and $\gamma$, and is previously introduced in Split LASSO (Cao et al., 2024a) in transformational sparsity problems. The effect of the hyperparameter $\nu$ for improving the selection power will be discussed by simulation experiments in Supplementary Material Section G.
With the above $W$-statistics, we are now ready to give the selectors of Model-X Split Knockoff. Let $q$ be our target FDR, two data-dependent threshold rules on a pre-set nominal level $q$ are defined as
@@
[…]
or $T_{q},T_{q}^{+}=+\infty$ if the respective set is empty, where $\mathcal{W}=\{|W_{j}|:j=1,2,\cdots,m\}\backslash\{0\}$. The selectors for the Model-X Split Knockoff (+) are respectively defined as
$\widehat{\mathcal{S}}=\{i:W_{i}\geq T_{q}\}\mathrm{\ or\ }\widehat{\mathcal{S}}=\{i:W_{i}\geq T_{q}^{+}\}.$
@@ 2.4 Connections with Model-X Knockoff
The Model-X Split Knockoff is a generalization of the Model-X Knockoff method into the problem of transformational sparsity, whose application includes but not limited to pairwise comparisons, wavelet transforms, fused LASSO, and trend filtering. In the following, we argue that the Model-X Split Knockoff can achieve better selection power compared with the Model-X Knockoff in the case where both methods are applicable.
@@
Consider the case where $D=I_{p}$ in the model (1), there exists a pair of $(A,\widetilde{A})$ in construction of the Model-X Split Knockoff, such that the selection power of such a construction achieves at least the same selection power as the Model-X Knockoff.
Proof.
Take $A=X$ in the construction of the Model-X Split Knockoff, then $\widetilde{X}^{A}=X-AI_{p}=0$. Under this condition, Model (3) and the exchangeability condition (4) collapse precisely into the Model-X Knockoff framework. While this particular construction therefore achieves equivalent selection power, alternative constructions may offer superior performance. ∎
@@ 3 FDR Control of Model-X Split Knockoff
In this section, theoretical guarantees for FDR control under the proposed Model-X Split Knockoff framework are provided. We start from the case where the distribution of $X$ is known and establish the exact FDR control. Then we apply the “leave one out” technique introduced in Barber et al. (2020) to establish robust FDR control when the distribution of $X$ can be estimated with a bounded deviation on KL-divergence.
3.1 Exact Model-X Split Knockoff
Now we are ready to present the false discovery rate control theorem of Model-X Split Knockoff when the distribution $X$ is known or bootstrapped exactly. In this setting, the set of random vectors $(\widetilde{X}^{A},A,\widetilde{A})$ exactly satisfies the exchangeability (4) and exact false discovery rate control is achieved.
Theorem 3.1 (FDR Control of Exact Model-X Split Knockoffs).
For any preset nominal FDR level $10<q<$ 1, if the set of random vectors $(\widetilde{X}^{A},A,\widetilde{A})$ satisfies the excangeability (4), the following statements holds for Model-X Split Knockoffs:
@@ 3.1 Exact Model-X Split Knockoff
    for the Model-X Split Knockoff, there holds
    $\mathrm{mFDR}=\mathbb{E}\left(\frac{|\{i\in\widehat{\mathcal{S}}\cap\mathcal{H}_{0}^{\gamma}\}|}{|\widehat{\mathcal{S}}|+1/q}\right)\leq q,$ (36)
@@
    for the Model-X Split Knockoff+, there holds
    $\mathrm{FDR}=\mathbb{E}\left(\frac{|\{i\in\widehat{\mathcal{S}}\cap\mathcal{H}_{0}^{\gamma}\}|}{|\widehat{\mathcal{S}}|\lor 1}\right)\leq q,$ (37)
where the $\mathrm{mFDR}$ represents the modified false discovery rate as introduced in Barber and Candès (2015) which adds $1/q$ in the denominator and should not have much impact when $|\widehat{\mathcal{S}}|$ is large.
Remark.
Theorem 3.1 will be a direct corollary of the robust FDR control Theorem 3.2 by setting the estimation error of $P_{X}^{*}$ to 0. The details will be discussed in Theorem 3.2, and the proof of Theorem 3.1 will therefore be omitted.
Example: Pairwise Comparisons.
Although the requirement that $(\widetilde{X}^{A},A,\widetilde{A})$ satisfies the excangeability (4) is generally unachievable in practice, there are some situations where such a condition reasonably holds.
For instance, consider the canonical example of pairwise comparisons in Section 2.2.2. In that case, the set of random vectors $(\widetilde{X}^{A}_{r},A,\widetilde{A})$ constructed by the bootstrap+ construction is based on bootstrapping (augmented with zeros) precisely the empirical distribution $P_{\mathbf{X}}$ of the random vector $X$. Therefore, it is straightforward to verify that $(\widetilde{X}^{A}_{r},A,\widetilde{A})$ satisfies the exchangeability (4), and exact false discovery rate control is achieved without estimating the distribution of $X$.
3.2 Approximate Model-X Split Knockoff
In practice, there are indeed many scenarios where the distribution of $X$ is unknown and needs to be estimated. In such cases, the set of random vectors $(\widetilde{X}^{A},A,\widetilde{A})$ only satisfies the exchangeability (4) approximately. To deal with such cases, we apply the “leave one out” technique introduced in Barber et al. (2020) to show that our method is robust to the estimation error on the distribution of $X$.
[…]
where $\widetilde{\mathbf{X}}^{A}_{i}$, $\mathbf{A}_{i}$, $\widetilde{\mathbf{A}}_{i}$ represent the $i$-th row of $\widetilde{\mathbf{X}}^{A}$, $\mathbf{A}$, $\widetilde{\mathbf{A}}$ receptively, and $\mathbf{A}_{i}(j)$ represents the vector obtained by substituting the $j$-th element of $\mathbf{A}_{i}$ with the $j$-th element of $\widetilde{\mathbf{A}}_{i}$.
Clearly, if $P_{X}=P^{*}_{X}$, e.g. the distribution of $X$ is known, the KL divergence $\widehat{\mathrm{KL}}_{j}=0$ for $j\in\{1,2,\cdots,m\}$. Meanwhile, in the cases where the true distribution $P^{*}_{X}$ is unknown and is estimated by $P_{X}$, $\widehat{\mathrm{KL}}_{j}$ measures the impact of the estimation error to the $j$-th feature based on the observations $\widetilde{\mathbf{X}}^{A},\mathbf{A},\widetilde{\mathbf{A}}$.
Now we are ready to state the main theorem on the robust false discovery control of the approximate Model-X Split Knockoffs.
Theorem 3.2 (Robust FDR control of Approximate Model-X Split Knockoffs).
Suppose the set of random vectors $(\widetilde{X}^{A},A,\widetilde{A})$ satisfy the exchangeability (4) when $X\sim P_{X}$. Then when $X\sim P^{*}_{X}$, for any $10<q<$ 1 and $\epsilon\geq 0$, there holds:
  1. 1.
    for the Model-X Split Knockoff, there holds
    $\mathbb{E}\left(\frac{|\{j:j\in\widehat{\mathcal{S}}\cap\mathcal{H}_{0}^{\gamma}\text{ and }\widehat{\mathrm{KL}}_{j}\leq\epsilon\}|}{|\widehat{\mathcal{S}}|+q^{-1}}\right)\leq q\cdot e^{\epsilon},$
    which leads to the following bound for modified FDR,
    $\mathrm{mFDR}=\mathbb{E}\left(\frac{|\{j:j\in\widehat{\mathcal{S}}\cap\mathcal{H}_{0}^{\gamma}\}|}{|\widehat{\mathcal{S}}|+q^{-1}}\right)\leq\min_{\epsilon\geq 0}\left\{q\cdot e^{\epsilon}+\mathbb{P}\left(\max_{j\in\mathcal{H}_{0}^{\gamma}}\widehat{\mathrm{KL}}_{j}>\epsilon\right)\right\},$
  2. 2.
    for the Model-X Split Knockoff+, there holds
    $\mathbb{E}\left(\frac{|\{j:j\in\widehat{\mathcal{S}}\cap\mathcal{H}_{0}^{\gamma}\text{ and }\widehat{\mathrm{KL}}_{j}\leq\epsilon\}|}{|\widehat{\mathcal{S}}|\vee 1}\right)\leq q\cdot e^{\epsilon},$
    which leads to the following bound for FDR,
@@ 3.2 Approximate Model-X Split Knockoff
Remark.
For the case that $P_{X}=P^{*}_{X}$, taking $\epsilon=0$ in Theorem 3.2 directly leads to Theorem 3.1. In this sense Theorem 3.1 can be regarded as a corollary of Theorem 3.2.
Theorem 3.2 states that the loss in the FDR control vanishes when the estimation on the distribution of $X$ becomes accurate. In the following, we will give a concrete example example on the robust FDR control when the distribution of $X$ is a normal distribution.
Example: Normal Distributions
In this following, we provide concrete computation of $\max_{j}\widehat{\mathrm{KL}}_{j}$ when $X$ follows the normal distribution $\mathcal{N}(\mathbf{0}_{p},\Sigma_{X})$. In this example, we construct the Model-X Split Knockoff in the way presented by Equation (25) and (30).
@@
Denote $\Theta:=\Sigma^{-1}_{X}$ and $\widehat{\Theta}:=\widehat{\Sigma}^{-1}_{X}$, where $\widehat{\Sigma}_{X}$ is an estimation of $\Sigma_{X}$. Suppose $X\sim\mathcal{N}(\mathbf{0}_{p},\Sigma_{X})$. Let $\Lambda_{X}$ the maximum eigenvalue of $\Sigma_{X}$
@@
for any choice of $\alpha>0$ such that $\widehat{\Sigma}_{X}-\alpha D^{T}D\succeq 0$ and $A\sim\mathcal{N}(0,\alpha I_{m})$. If $\frac{\log{(m+p)}}{n}=o(1)$, with probability at least $1-\frac{2}{p}-\frac{2}{m}$, there holds
@@
In many scenarios including total variation or multiple comparisons, $\|D_{j}\|_{2}$ is at constant levels, so we only need to focus on the estimation error of $\Theta$ to establish bounds on $\delta_{\Theta}$ and $\max_{j}\widehat{\mathrm{KL}}_{j}$. To be specific, the estimation error $\|\Theta-\widehat{\Theta}\|_{2}$ needs to be $o_{p}\left(\frac{1}{\sqrt{n\log p}}\right)$ in order to make $\max_{j}\widehat{\mathrm{KL}}_{j}$ sufficiently small.
To achieve this goal, unlabeled data can be utilized for estimation of $\Theta$, which are typically more accessible and thus considerably outnumber labeled data. To be specific, consider the case where there are $N\gg n$ unlabeled samples. Given that $X$ follows a multivariate Gaussian distribution, a simple sample covariance matrix $S_{N}:=\frac{1}{N}(X^{u})^{\top}X^{u}$ computed over unlabeled data $X^{u}$ can be used as the estimation, which has $\|\Theta-\widehat{\Theta}\|_{2}\sim O_{p}\left(\sqrt{\frac{p\log{p}}{N}}\right)$. In this case, $\max_{j}\widehat{\mathrm{KL}}_{j}\sim o_{p}(1)$ is obtained as long as
$2\delta_{\Theta}\sqrt{n\log{m}}\asymp\sqrt{\frac{p\log{p}}{N}}\sqrt{pn\log{m}}=o_{p}(1),$
i.e. if $N\gg np^{2}\log^{2}{(m+p)}$. Further, if the inverse covariance matrix is sparse, one can employ the Graphical Lasso for estimation:
$\widehat{\Theta}_{\lambda}=\arg\min_{\Theta\succeq 0}\left\{-\log\operatorname{det}(\Theta)+\left\langle\Theta,S_{N}\right\rangle+\lambda\sum_{j\neq k}\left|\Theta_{jk}\right|\right\}.$
Under certain conditions, Ravikumar et al. (2011) shows that $\|\Theta-\widehat{\Theta}\|_{2}\asymp\sqrt{\frac{d^{2}\log{p}}{N}}$, where $d$ denotes the maximum degree. This property only requires $N\gg npd^{2}\log^{2}{p}$ to make $\max_{j}\widehat{\mathrm{KL}}_{j}$ to be $o_{p}(1)$.

### v2

@@ 2.2.2 Construction for Pairwise Comparisons
In this section, we discuss the case of pairwise comparisons where the distribution of the design matrix $X$ can not be approximated by normal distributions.
For example, let the pairwise comparisons be represented by a design matrix $\mathbf{X}\in\mathbb{R}^{n\times p}$ and responses $\mathbf{Y}\in\mathbb{R}^{n}$. Each row $X\in\mathbb{R}^{p}$ of $\mathbf{X}$ encodes a single comparison between a random pair of items $(j,k)$ where $j<k$, such that its entries are defined as $X_{j}=1$, $X_{k}=-1$ and zeros elsewhere. Each element $Y\in\mathbb{R}$ of $\mathbf{Y}$ is modeled from
@@
where $\beta^{*}\in\mathbb{R}^{p}$ represents the unknown true scores for $p$ objects, and $f(x)$ is a cumulative distribution function, symmetric at 0, i.e. f(x)=1f(x)f(x)=1-f(-x). To conduct pairwise comparisons, let $D\in\mathbb{R}^{\frac{p(p-1)}{2}\times p}$ be the graph difference operator of a fully connected graph. [^2] The construction detailed in this section is applicable when $D$ contains all the edges present in $\mathbf{X}$. Should this condition not be met, one can simply augment $D$ to include all required edges. Then pairs with large differences will be selected in model (1).
[…]
Bootstrap+ Constructions. Now we will present a novel method of generating Split Knockoff copies, initiated by an augmented bootstrap resampling procedure, so we term it as the bootstrap+ construction. First, each row $(X^{r},Y^{r})$ of the new design matrix $\mathbf{X}^{r}$ and response vector $\mathbf{Y}^{r}$ is a bootstrap sampling from the distribution of $(X,Y)$ and augmented by zeros by:
@@
    $\mathbb{P}(X^{r}=0_{p})=\mathbb{P}(X^{r}\neq 0_{p})=\frac{1}{2}$.  [^3] One can not use a general probability $\mathbb{P}(X^{r}=0_{p})=p_{0}\neq 1/2$ here since $p_{0}\neq 1/2$ will destroy the exchangeability between $\tilde{A}$ and $A$.
@@
    When $X^{r}\neq 0_{p}$, $(X^{r},Y^{r})$ takes a random row from $(\mathbf{X},\mathbf{Y})$ respectively.
This resampling enables the following batch construction of the Split Knockoff copies based on $(X^{r},Y^{r})$ at a much lower computational cost.
@@
    The random vector $A\in\mathbb{R}^{m}$, where $m=\frac{p(p-1)}{2}$, is generated to satisfy $\widetilde{X}^{A}_{r}:=X^{r}-AD=0_{p}$, i.e. if $X^{r}=0_{p}$ then $A=0_{m}$; if $X^{r}$ takes a random row $i$ of $\mathbf{X}$ then $A=e_{i}$ selects the same row in $D$, where $\{e_{i}\}$ is the canonical basis of $\mathbb{R}^{m}$.
@@
    The random vector $\widetilde{A}$ is generated based on $A$ by complementarily selecting zero when $A$ is non-zero, i.e. $\mathbb{P}(\widetilde{A}=0_{m}|A\neq 0_{m})=1$, or uniformly from any canonical basis vector $e_{i}$ when $A$ is zero, i.e. $\mathbb{P}(\widetilde{A}=e_{i}|A=0_{m})=\mathbb{P}(A=e_{i}|A\neq 0_{m})$.
With the above construction, one can verify that $\widetilde{X}^{A}_{r}\equiv 0_{p}$, while $A,\widetilde{A}]\in\mathbb{R}^{2m}$ is the one-hot vector with one element being 1 and other elements being zeros. Moreover, $\mathbb{P}(A_{j}=1)=\mathbb{P}(\widetilde{A}_{j}=1)$ for all $j$, which leads to the exchangeability.
This procedure clearly lets $\widetilde{X}^{A}_{r}$ independent from $A$, which could help improve the selection power as demonstrated in the Normal distribution case. Simulation experiments in Section 4.2 validate the selection power improvement with less computational cost compared with the sequential construction. Moreover, the bootstrap+ construction only requires evaluation of $\mathbb{P}(A=e_{i}|A\neq 0_{m})$, which can be directly computed from the known empirical distribution $P_{\mathbf{X}}$ of the random vector $X$. This leads to exact FDR control to be discussed later in Section 3.
@@ 2.3 Computation of $W$-statistics
With the above construction, we are now ready to compute the routine $W$-statistics
$W_{i}=w_{i}\left((\widetilde{X}^{A},A,\widetilde{A}),Y\right),$ (18)
for $i\in\{1,2,\cdots,m\}$, similar to (Barber and Candès, 2015, Candès et al., 2018). Similarly, the $W$-statistics needs to satisfy the flip-sign property:
$w_{i}\left((\widetilde{X}^{A},A,\widetilde{A})_{\text{swap}(S)},Y\right)=\begin{cases}w_{i}\left((\widetilde{X}^{A},A,\widetilde{A}),Y\right),&i\notin S,\\ -w_{i}\left((\widetilde{X}^{A},A,\widetilde{A}),Y\right),&i\in S,\end{cases}$ (19)
for any $S\subseteq\{1,2,\cdots,m\}$ and any $i\in\{1,2,\cdots,m\}$.
There are many approaches, as shown in (Barber and Candès, 2015, Candès et al., 2018), in constructing such $W$-statistics satisfying the above property. In this paper, we adopt the following specific construction proposed by (Cao et al., 2024a) that searches in the relaxed neighborhood of the subspace constraint to optimize the selection power. For hyperparameters $\lambda>0$, $\nu>0$, we optimize:
@@
where the loss function $L$ is related to the model or link function, e.g., mean squared loss for linear regression, and cross-entropy loss for logistic regression. The $W$-statistics is constructed as $W=|\gamma(\lambda)|-|\widetilde{\gamma}(\lambda)|$.
The hyperparameters $\lambda$ and $\nu$ can be optimized to maximize the selection power. Here, $\lambda$ is the well-known $\ell_{1}$ regularization parameter. Meanwhile, $\nu$ is the parameter which controls the Euclidean relaxation gap between $D\beta$ and $\gamma$, and is previously introduced in Split LASSO (Cao et al., 2024a) in transformational sparsity problems. The effect of the hyperparameter $\nu$ for improving the selection power will be discussed by simulation experiments in Supplementary Material Section G.
With the above $W$-statistics, we are now ready to give the selectors of Model-X Split Knockoff. Let $q>0$ be our target FDR, two data-dependent threshold rules on a pre-set nominal level $q$ are defined as
$\bullet\mbox{(Model-X Split Knockoff)}\ \ \ T_{q}=\min\left\{\lambda\in\mathcal{W}:\frac{|\{i:W_{i}\leq-\lambda\}|}{1\vee|\{i:W_{i}\geq\lambda\}|}\leq q\right\},$
$\bullet\mbox{(Model-X Split Knockoff+)}\ \ \ T_{q}^{+}=\min\left\{\lambda\in\mathcal{W}:\frac{1+|\{i:W_{i}\leq-\lambda\}|}{1\vee|\{i:W_{i}\geq\lambda\}|}\leq q\right\},$
@@ 2.4 Connections with Model-X Knockoff
The Model-X Split Knockoff is a generalization of the Model-X Knockoff method into the problem of transformational sparsity, whose application includes, but is not limited to pairwise comparisons, wavelet transforms, fused LASSO, and trend filtering. Next, we argue that the Model-X Split Knockoff can achieve better selection power compared with the Model-X Knockoff when both are applicable.
@@
Consider the case $D=I_{p}$ in the model (1), there exists a pair of Model-X Split Knockoff copies (A,A~)($\widetilde{A}$), such that Model-X Split Knockoff with (A,A~)($\widetilde{A}$) achieves at least the same selection power as the Model-X Knockoff.
The proof of this proposition is straightforward. Take $A=X$ in the construction of the Model-X Split Knockoff, then $\widetilde{X}^{A}=X-AI_{p}=0$. Under this condition, the model (3) and the exchangeability condition (4) degenerate precisely into the Model-X Knockoff framework. Therefore, Model-X Split Knockoff with this choice of $A$ have the same selection power with the Model-X Knockoff.
@@ 3 FDR Control of Model-X Split Knockoff
In this section, we provide the robust FDR control guarantees under the proposed Model-X Split Knockoff framework. To be specific, we apply the “leave-one-out” technique introduced in Barber et al. (2020) to establish robust FDR control when the distribution of $X$ can be estimated with a bounded deviation on the KL-divergence.
In practice, there are many scenarios where the distribution of $X$ is unknown and needs to be estimated. To be specific, let $P_{X}$ be the estimated distribution [^4] Here, $P_{X}$ and $P^{*}_{X}$ represent the probability mass function for discrete random variables and probability density function for continuous random variables. of $X$, and $P^{*}_{X}$ be the true distribution of $X$. Consider the following sample Kullback–Leibler (KL) divergence for $j\in\{1,2,\cdots,m\}$:
where $\widetilde{\mathbf{X}}^{A}_{i}$, $\mathbf{A}_{i}$, $\widetilde{\mathbf{A}}_{i}$ represent the $i$-th row of $\widetilde{\mathbf{X}}^{A}$, $\mathbf{A}$, $\widetilde{\mathbf{A}}$ receptively, and $\mathbf{A}_{i}(j)$ represents the vector obtained by substituting the $j$-th element of $\mathbf{A}_{i}$ with the $j$-th element of $\widetilde{\mathbf{A}}_{i}$. Clearly, if $P_{X}=P^{*}_{X}$, e.g. the distribution of $X$ is known, the KL divergence $\widehat{\mathrm{KL}}_{j}=0$ for $j\in\{1,2,\cdots,m\}$. Meanwhile, in the cases where the true distribution $P^{*}_{X}$ is unknown and is estimated by $P_{X}$, $\widehat{\mathrm{KL}}_{j}$ measures the impact of the estimation error to the $j$-th feature based on the observations $\widetilde{\mathbf{X}}^{A},\mathbf{A},\widetilde{\mathbf{A}}$.
Now we are ready to state the main theorem on the robust false discovery control of the Model-X Split Knockoffs.
Theorem 3.1.
[…]
@@
    For the Model-X Split Knockoff, there holds $\mathbb{E}\left(\frac{|\{j:j\in\widehat{\mathcal{S}}\cap\mathcal{H}_{0}^{\gamma}\text{ and }\widehat{\mathrm{KL}}_{j}\leq\epsilon\}|}{|\widehat{\mathcal{S}}|+q^{-1}}\right)\leq q\cdot e^{\epsilon}$, leading to the following bound of $\mathrm{mFDR}$ defined in Barber and Candès (2015),
    $\mathrm{mFDR}=\mathbb{E}\left(\frac{|\{j:j\in\widehat{\mathcal{S}}\cap\mathcal{H}_{0}^{\gamma}\}|}{|\widehat{\mathcal{S}}|+q^{-1}}\right)\leq\min_{\epsilon\geq 0}\left\{q\cdot e^{\epsilon}+\mathbb{P}\left(\max_{j\in\mathcal{H}_{0}^{\gamma}}\widehat{\mathrm{KL}}_{j}>\epsilon\right)\right\}.$
@@
    For the Model-X Split Knockoff+, there holds $\mathbb{E}\left(\frac{|\{j:j\in\widehat{\mathcal{S}}\cap\mathcal{H}_{0}^{\gamma}\text{ and }\widehat{\mathrm{KL}}_{j}\leq\epsilon\}|}{|\widehat{\mathcal{S}}|\vee 1}\right)\leq q\cdot e^{\epsilon}$, which leads to the following bound of $\mathrm{FDR}$,
@@
Theorem 3.1 states that the loss in the FDR control vanishes when the estimation on the distribution of $X$ becomes accurate. For the case that $P_{X}=P^{*}_{X}$, taking $\epsilon=0$ in Theorem 3.1 leads to the exact FDR control. In the following, we will give one concrete example of such exact FDR control in pairwise comparisons, and one concrete example of robust FDR control when the distribution of $X$ is normal.
[…]
Robust FDR Control Example: Normal Distributions. In the following, we provide concrete computation of $\max_{j}\widehat{\mathrm{KL}}_{j}$ when $X$ follows the normal distribution $\mathcal{N}(\mathbf{0}_{p},\Sigma_{X})$. In this example, we construct the Model-X Split Knockoff as in (16).
@@
Denote $\Theta:=\Sigma^{-1}_{X}$ and $\widehat{\Theta}:=\widehat{\Sigma}^{-1}_{X}$, where $\widehat{\Sigma}_{X}$ is an estimation of $\Sigma_{X}$. Suppose $X\sim\mathcal{N}(\mathbf{0}_{p},\Sigma_{X})$. Let $\Lambda_{X}$ the maximum eigenvalue of $\Sigma_{X}$, and define
@@
for any choice of $\alpha>0$ such that $\widehat{\Sigma}_{X}-\alpha D^{\top}D\succeq 0$ and $A\sim\mathcal{N}(0,\alpha I_{m})$. If $\frac{\log{(m+p)}}{n}=o(1)$, with probability at least $1-\frac{2}{p}-\frac{2}{m}$, there holds
@@
In many scenarios, $\|D_{j}\|_{2}$ is at constant levels. Therefore, the estimation error $\|\Theta-\widehat{\Theta}\|_{2}$ being $o_{p}\left(\frac{1}{\sqrt{n\log p}}\right)$ is sufficient to make $\max_{j}\widehat{\mathrm{KL}}_{j}$ sufficiently small. Unlabeled data can help to achieve this goal with details being discussed in the remark of Supplementary Material Section E.